OpenStudy (anonymous):

What is the simplified fraction form of 15xsqrt(3)/sqrt(75x^2)?

OpenStudy (jhannybean):

\[\large \frac{15x\sqrt{3}}{\sqrt{75x^2}}\] first simplify the denominator \[\large \sqrt{75 x^2} = \sqrt{ \color{blue}{25}\cdot 3 \color{magenta}{\cdot x \cdot x}}\]

OpenStudy (jhannybean):

\[\large \sqrt{75 x^2} = \sqrt{ \color{blue}{25}\cdot 3 \color{magenta}{\cdot x \cdot x}}= \sqrt{\color{blue}{5 \cdot 5} \cdot 3 \cdot \color{magenta}{x \cdot x}}\] anything you can pull out of the square root?

OpenStudy (anonymous):

5

OpenStudy (jhannybean):

and what else? :)

OpenStudy (anonymous):

the x's

OpenStudy (jhannybean):

good job. that leaves you with \[\large \color{blue}5\color{magenta}x\sqrt{3}\] correct?

OpenStudy (anonymous):

Yes

OpenStudy (jhannybean):

Now that we've factored out our denominator,lets rewrite the problem :D \[\large \frac{15x\sqrt{3}}{5x\sqrt{3}}\] is there any like-terms we can cancel out here?

OpenStudy (anonymous):

the sqrt(3)

OpenStudy (jhannybean):

Yes that's one. Anything else?

OpenStudy (anonymous):

The x's

OpenStudy (jhannybean):

Thats two! anything we can simplify? :)

OpenStudy (anonymous):

Ya the 15 and 5 go to 3

OpenStudy (jhannybean):

\[\large \frac{15\cancel x\ \cancel{\sqrt{3}}}{5\cancel x\ \cancel{\sqrt{3}}}\]

OpenStudy (jhannybean):

good! and you got your answer :)

OpenStudy (anonymous):

So the answer is 3

OpenStudy (jhannybean):

Yep.

OpenStudy (jhannybean):

And to check this we can plug 3 back into our original equation \[\large \frac{15x\sqrt{3}}{\sqrt{75x^2}}\] x = 3 \[\large \frac{15(3)\sqrt{3}}{\sqrt{75(3)^2}} = 3\]

OpenStudy (jhannybean):

\[\large \frac{45\sqrt{3}}{\sqrt{675}}= 3\] the reduced form. plug this into your calculator :)